# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ III-A Localization and Mapping
GPS can directly provide the location of the vehicle, but the accuracy is affected by a list of factors, such as atmospheric conditions, satellite geometry, signal blockage (buildings, trees, etc.), multipath interference, and the quality of the GPS receiver. Differential GPS (DGPS) systems and augmentation techniques can be used to enhance GPS accuracy by correcting some of these error sources, but cannot fulfill the accuracy which is extremely important when it is small-scale cars. Typically, an IMU consists of various sensors, such as accelerometers and gyroscopes to measure and report acceleration and angular velocity in three dimensions: pitch, roll, and yaw with high rates. However, errors such as sensor drift and integration errors can accumulate over time with IMU, leading to inaccuracies in long-term measurements. To address the issues, the two sensors are often integrated through sensor fusion methods, such as Kalman filter (KF) [106] or its variations as extended Kalman filter (EKF) or unscented Kalman filter (UKF) [107], to enhance accuracy and reliability in determining the position and orientation. As in the studies by [108, 109, 110], that use BARC small-scale cars, KF makes the fusion of IMU, wheel encoders, and indoor GPS measurements to achieve an accurate localization. In recent research, [111] presents an approach to use KF to fuse multiple-camera images with the AprilTag system and wheel odometry. The results are validated with the OptiTrack Motion Capture System (MoCap). OptiTrack system can also be used as a global localization system and provide high accuracy pose estimation but with higher cost [112], more detail will be discussed in Section IV-E.
@@
For small-scale vehicles, the primary sensor for localization and mapping components utilized is the LiDAR. They collaborate with a pre-defined map or SLAM technology to navigate and position the vehicle effectively within its environment [113]. A well-established method for localizing robot cars with LiDAR sensors on the pre-defined map involves the use of a particle filter (PF) as explored in [45] with F1TENTH. However, the computational demands of the PF present challenges, especially for computation resources and space-constrained small-scale robot cars. To enhance performance, [114] introduces the Compressed Directional Distance Transform (CDDT) on the RACECAR platform. This method focuses on expediting ray casting within 2D occupancy grid maps and accelerating sensor model computations to mitigate computational expenses. While comprehensive survey papers delve into various SLAM techniques [115, 116, 117, 118], encompassing feature-based, LiDAR-based, visual, and graph SLAM, this study offers a concise overview of three prevalent SLAM methods employed with small-scale cars: GMapping [119, 120, 121], HectorSLAM [122, 123, 124, 125, 72], and Cartographer [126, 127, 128].
GMapping, utilizing a particle filter-based approach, stands as one of the most widely used SLAM algorithms, adept at handling non-linear sensors and non-Gaussian noise [129]. To address computational expenses inherent in particle filters, the Rao-Blackwellized Particle Filters (RBPF) algorithm is introduced to combine odometry with scan-matching techniques. Based on particles representing a sampling from a multimodal map proposal distribution, this approach effectively closes loops without necessitating additional methods. The efficacy of GMapping hinges on both 2D LiDAR data and odometer information, utilizing the latter significantly to curtail the number of particles involved in the mapping process. In contrast, HectorSLAM exclusively depends on LiDAR sensors and leverages the high sampling frequency of modern 2D LiDAR [130]. It swiftly provides a 2D pose estimate at the scan rate utilizing a scan-matching algorithm [130]. Upon receiving raw LiDAR data and joint values, HectorSLAM undergoes preprocessing before importing the data to the scan matching module, applying the Gauss-Newton approach [131] to estimate the optimal pose. As robots traverse longer trajectories and handle substantial data volumes, the limitations of particle-based approaches become evident. Graph optimization algorithms emerge as more reliable and efficient solutions for managing these large-scale datasets. As a typical Graph SLAM, the Cartographer takes odometry data as an initial estimate for the local optimizer [132]. This optimizer focuses on determining the most probable relative pose between the scan and local submap coordinate frames. These relative poses serve as constraints for a Sparse Pose Adjustment optimizer, specifically engineered to address loop closure challenges. Cartographer operates as a dual-pronged approach to the SLAM problem, comprising the front-end (Local SLAM) and back-end (Global SLAM). The front-end executes critical tasks like scan matching for loop closure detection and estimating the pose. Conversely, the back-end optimizes poses regularly by forming a graph to minimize loop-closing constraints, thereby minimizing successive pose errors.
@@
Moreover, a compelling approach gaining traction involves leveraging ML methods to address mapping and localization challenges. One notable instance is the GALNet proposed in [133] implemented on the Autominy platform. GALNet employs a Deep Neural Network (DNN), utilizing two timestamps of inertial, kinematic, and wheel velocity data to estimate poses effectively. With the success of Transformer architecture in various research fields, [134] introduces the Perception-Action Causal Transformer (PACT) architecture. This model constructs a representation from sensor data by autonomously predicting states and actions over time, laying the groundwork for subsequent task-specific networks for localization and mapping. Pre-trained models are employed initially and later fine-tuned for specific tasks, demonstrating validation on MuSHR with LiDAR sensor data. For image-based localization, [135] proposes an effective framework leveraging a pre-trained local feature transformer (LoFTR). This framework employs a constrained 3D projective transformation between consecutive key images to establish a visual map on Jetbot. These studies underscore the growing interest and potential of ML methodologies, including neural networks and transformer architectures, in revolutionizing mapping and localization paradigms by offering robust, data-driven solutions that outperform traditional approaches.
III-B Path planning and following
Similar to localization and mapping, path planning and path following can be regarded as distinct tasks for small-scale cars. While often studied together, both are essential components that are merged to form a cohesive navigation process within autonomous systems.
III-B1 Path planning
Path planning in autonomous driving refers to the process by which a self-driving vehicle determines a safe and efficient route from its current location to a desired destination while navigating through its environment [136]. This task involves creating a trajectory or path that the vehicle can follow, considering various factors such as obstacles, road conditions, and traffic regulations, and aims to ensure safe and reliable navigation while optimizing factors like travel time, and energy efficiency. Path planning algorithms vary in complexity and we specify them into traditional [137, 138, 139] and ML-based techniques [140, 141].
@@ III-B1 Path planning
The most traditional method is Dijkstra’s algorithm [137], which is used in [142] for path planning with SLAM methods to obtain the global map information. Among many applications of the A* [138], [40] uses the A* algorithm to obtain optimal paths for a racing car, and [143] integrates the Hybrid A* algorithm [144] with a learned model utilizing a cost function. This integration aimed to tackle unforeseen obstacles encountered during navigation, which often escape the detection of traditional global path planning algorithms. In the work by [145] and [146], genetic algorithms (GA) are investigated to generate a global path for a mobile robot in the grid-based environment. Furthermore, to extend the planning capabilities beyond the sensor horizon, [147] utilizes a generative neural network trained on real-world data to predict occupancy maps beyond sensor limitations, see Fig. 4. This predictive capability assists in facilitating planning processes. Authors leverage the Rapidly-Exploring Random Tree (RRT) algorithm [139] to generate a global path, followed by a local planner that orchestrates trajectories until the end of the predicted map.
Figure 4: Overall perception and control pipeline in [147]. The perception module receives onboard sensor data and produces a predicted occupancy map using a U-Net style generative neural network. The control algorithm receives the robot state, predicted occupancy map, and goal point and generates collision-free trajectories with RRT.
@@
Path planning, treated as an optimization problem, has also seen the successful application of various ML methods, resulting in notably high-performance outcomes. For instance, [148] uses Q-Learning, a classical RL algorithm, to generate a global path for robots, which achieves a shorter and smoother path compared with the RRT algorithm. A framework is introduced in [149] comprising a mapper, global policy, and local policy for image-based navigation for Jetbot. The mapper undergoes supervised learning using camera images to generate an occupancy grid map. Subsequently, a global policy employing DRL techniques takes both the map and the position as input to determine the long-term goal. Following this, a planner, employing the A* algorithm, computes the short-term goal, which is then forwarded to the local planner for further execution.
Even with prior maps and acquired information, managing uncertainty during planning and driving remains a significant challenge. To address these uncertainties, [150] introduces a model-based RL algorithm incorporating a probabilistic dynamic model. This method aims to mitigate uncertainty during planning stages and prevent shortsighted decisions, illustrated in Fig. 5. In the related approach, [151] uses Bayesian Residual Policy Optimization (BRPO) with an ensemble of expert policies. Authors train the policy ensemble with BRPO to diminish overall system uncertainty, enabling safe navigation within partially observable environments. Demonstrated on the MuSHR platform, proposed framework integrates a global localization system, ensuring destination-reaching capabilities while avoiding collisions with other moving vehicles.
Figure 5: An overview of the method in [150]. The probabilistic dynamics model is trained using a multistep loss that considers how uncertainty propagates. The trajectory tracking controller is used to predict a distribution of closed-loop trajectories. The divergence constrained optimizer is used to find a closed-loop trajectory with low divergence for MuSHR cars.
III-B2 Path following
Once the path is established, the path following becomes the subsequent step, encompassing the capability of the vehicle to precisely track the designated path in real time. Here, the control system interprets the planned trajectory and orchestrates the essential actions required for the vehicle to adhere to the intended path. Path following necessitates a continual adjustment of the movements based on real-time sensor feedback and environmental alterations, ensuring the vehicle maintains the desired trajectory accurately. Like path planning, solutions for path following tasks encompass a wide spectrum, including classical control methods like the PID controller, kinematic controller [152, 153, 154], MPC controller [155, 156, 157, 158], and extending to ML methods [159, 160, 149, 161].
In [154], an adaptive trajectory tracking control scheme, as shown in Fig. 6, is introduced with adjustable gains to facilitate adherence to predefined paths. The designed adaptive control gains aimed to streamline tuning efforts, augment the convergence rate of tracking errors, and elevate trajectory tracing performance. Meanwhile, the strategy proposed by [158] involves expert interventions with pre-existing LiDAR-derived maps to further improve performance. As for the ML methods, in the work presented by [159], error states relative to reference trajectories are determined using vehicle state data obtained from sensor fusion involving IMU and GPS with an EKF. A Neural Network (NN) is trained using IL methods with a dataset collected from human-controlled and MPC-controlled driving scenarios. Trained NN maps error states to control commands. To avoid human-induced errors and large datasets, RL has also been widely adopted in path following tasks, as evidenced by [160], where the Deep Deterministic Policy Gradient (DDPG) algorithm is applied to guide an Autominy car along the desired path, aiming to minimize cross-track errors. Similarly, in the study by [149], a global policy is determined using DRL, while a local planner processes received images with another DRL agent to derive the final action for the path following. To address the data inefficiency challenges with RL, [161] then introduces a policy gradient-based policy optimization framework leveraging a first-principles model for path following task with a JetRacer. This framework facilitated the learning of precise control policies with limited real-world data.
Figure 6: Architecture of the proposed adaptive trajectory tracking control system in [154].
III-C Lane Keeping
Next, we discuss another fundamental driving behavior for small-scale vehicles, lane keeping, which has been a pivotal focus in extensive research across various platforms [7, 10, 11]. While the lane keeping task shares similarities with the path following, there are distinct differences between the two. Path following involves adhering to a predetermined path and utilizing real-time feedback to maintain desired trajectories safely. On the other hand, the objective of the lane keeping task is to guide the small-scale vehicle within the designated lane, conforming to specific tolerances for deviation from the lane center. This task requires the small-scale car to implicitly follow a path dynamically defined by the vehicle in real time. Consequently, this task predominantly relies on visual input, utilizing RGB cameras to monitor the position within the lane continuously. According to the methods used, we categorize research works into two groups, traditional methods and ML-based methods.
III-C1 Traditional methods
Traditional methodology typically comprises two integral components: perception and control, similar to various other driving tasks. The perception module is responsible for processing sensor data, often obtained from cameras, to extract crucial information necessary for lane keeping. It can be divided into two parts, lane detection for analyzing the sensor data to identify and track lane markings on the road; and feature extraction for calculating relevant metrics such as lateral deviation and orientation deviation to determine the position of the car within the lane. The control module then utilizes the information extracted from the perception module to regulate the trajectory and ensure it remains within the designated lane. Control algorithms such as PID controllers or other rule-based systems are employed to interpret the deviation data and adjust the steering angle accordingly.
In [7], a multi-step image processing pipeline, shown in Fig. 7, is devised to address the lane keeping task for Dukietown using traditional computer vision methodologies. With k-means clustering, Canny filter [162], Hue-Saturation-Value (HSV) colorspace thresholding, and probabilistic Hough transform [163], individual line segments are extracted, afterwards, A nonlinear non-parametric histogram filter is employed to estimate the lateral displacement concerning the right lane center and the angle offset relative to the center axis. These estimates hold immense significance within the lane-keeping task, serving as critical inputs for the subsequent control module. The control module employs a PID controller, utilizing these estimated parameters in real time to minimize deviations and steer the vehicle along optimal trajectories within the designated lane. Later [164] presents another similar pipeline for the Donkeycar platform, in which the H layer of HSV colored image is used to go through Gaussian kernel filter for noise filtering and Sobel edge detector [165], following the probabilistic Hough transform for lane detection. Then a list of rules is proposed to compute the steering angle for the car based on the detected lines. While traditional methods serve well for basic detection and control tasks, their limitations in generalization prompt the need for more advanced methodologies.
Figure 7: Different outputs of the multi-step image processing pipeline proposed by [7] and later employed in [94, 95]. It starts from the image input from the camera, proceeds to detect road markings according to the color with HSV thresholding, detects marking edges with canny filter, and finalizes with detection marking used to estimate the lateral displacement and angle offset.
III-C2 ML-based methods
Similar to the other driving tasks, ML-based methods are used to improve lane keeping capabilities and their generalization. Here, we divide them into two subcategories, namely modular system and end-to-end system, as in Fig. 2.
Modular system
For the lane keeping task, an ML-based modular system typically replaces traditional perception or control systems with more sophisticated ML-based algorithms. In [166], Convolutional Neural Network (CNN) models are used to predict the lateral displacement and the angle offset. These predictions were then used in conjunction with the lane filter and PID controller derived from [7] to control the Duckiebot. Besides replacing the perception module, [95, 94] use the same perception module from [7], but replaces the traditional controller with DRL algorithms that use the outputs from the perception module and achieves better performance than simple PID controller with multitask driving, not only lane keeping. With this modular system, the Sim2Real capabilities of the Duckiebot are also improved. For the Sim2Real capabilities improvement or generalization, [43] presents a modular system specifically tailored for lane keeping tasks. First, the perceptual system is trained with a mixture of simulation and real world datasets and map images into a shared latent space. Then, a list of DRL algorithms is employed, utilizing these latent features as inputs to control the Donkeycar. Furthermore, the evolution of transformer architectures, particularly the Vision Transformer (ViT) [167], has gathered increasing interest, competing with conventional CNN. For lane keeping tasks, for instance, in the work by [168], a pre-trained ViT is employed as a backbone network, see illustration in Fig. 8. This ViT model is further fine-tuned using limited datasets obtained from the driving environment to generate segmentation of the environment. Subsequently, a potential-field-based controller utilizes the segmentation output generated by the ViT as input and produces steering commands for control.
Figure 8: In [168], image patches are encoded using a ViT and predict a class label for each patch. Then, the coarse segmentation output is used for a potential-field based controller.
End-to-end system
A more highly researched system is the end-to-end system for lane keeping tasks, in which the agent receives raw sensor input and output control commands for lane keeping tasks. This control framework encompasses both IL and DRL. IL involves the system learning from expert demonstrations or human-driven data to emulate behaviors, decision-making, and actions observed in provided training demonstrations. Common IL algorithms include Behavioral Cloning (BC) [169], Generative Adversarial Imitation Learning (GAIL) [170], Inverse Reinforcement Learning (IRL) [171], Dataset Aggregation (DAgger) [172], and others. Since BC is a straightforward approach in IL that involves training a model to mimic expert behavior by directly mapping observed states to corresponding actions without understanding the underlying decision-making process, it is one of the foundational algorithms used for lane keeping tasks. For example, in [49, 173], CNNs are trained with image data to predict directly the steering angle for control under various evaluation metrics within the DeepRacer platform. However, the most crucial drawback of BC is the compounding errors, especially when the model encounters states not seen during training, leading to divergence from expert behavior. It is sensitive to errors in the training data and might not capture the nuanced decision-making aspects of the expert. To improve the generalization and transferability of Sim2Real in the Donkeycar platform, [92] proposes incorporating uncertainty during training. Similarly, [174] augments the dataset by image style transfer [175]. BC, GAIL, and DAgger are compared by [176] within the Duckietown environment, and apply Domain Adaption to facilitate the Sim2Real transfer issue. In addition to IL, DRL stands out as another commonly used method with raw sensor data as input. [10, 177, 178] employ raw images directly into DRL controllers, while to speed up convergence, [179, 180, 99] preprocess input images, employing simple techniques to reduce input dimensions. For enhanced control smoothness, [181] combines Conditioning for Action Policy Smoothness (CAPS) with RL, addressing jerky control issues. For Sim2Real transfer capabilities, [38] includes delays and sampling rate as additional agent observations at training time to improve the robustness of DRL policies for Sim2Real transfer of the DeepRacer platform. Additionally, besides image inputs, raw LiDAR data serves as input in studies such as [47], where DRL algorithms navigate an F1TENTH car within structured environments using raw LiDAR input.
To further enhance convergence speed and boost generalization capabilities, the end-to-end systems often integrate autoencoders [182]. It consists of an encoder and a decoder: the encoder compresses input data into a lower-dimensional representation through a sequence of neural network layers, while the decoder reconstructs the original input from the compressed representation in the latent space. The training objective of an autoencoder is to minimize the difference between input data and the reconstructed output. This encourages the network to learn an efficient representation of the data in the latent space, subsequently utilized as input for control networks. As outlined in [183, 184], autoencoders are initially trained to compress image data into lower-dimensional representations. Following this compression, BC is employed to train Duckiebots using expert trajectories. Autoencoders are frequently integrated with DRL algorithms, for example, [42, 185] utilize variational autoencoders (VAE) [186] to compress images and use the resulting latent features as inputs for DRL algorithms. This approach, compared to DRL algorithms with raw image inputs, demonstrates faster convergence and improved learning efficiency.
III-D Car following and overtaking
Among the previously mentioned tasks, car following and overtaking with small-scale cars remain relatively understudied areas. Nonetheless, they hold significant importance for research on autonomous small-scale vehicles.
III-D1 Car following
Car following maneuver is the capability of an autonomous vehicle to maintain a safe and appropriate distance from the vehicle ahead while driving on the road. It requires the vehicle to track and follow the movements of the leading vehicle, adjusting its speed and position accordingly to ensure a safe, comfortable, and efficient driving experience. The existing literature on car following maneuvers with small-scale cars is relatively scarce, with a few notable studies standing out. A DDPG algorithm with an extended look-ahead approach is proposed in [187] for longitudinal and lateral control within vehicle platooning scenarios. For perception, LiDAR technology gauges inter-vehicle distance, while an IMU provides acceleration data. Additionally, a V2V system employing Wi-Fi communication transmits leader information to the follower vehicle. Similarly, in [188], a Cooperative Adaptive Cruise Control (CACC) system is implemented utilizing Deep Q-learning (DQN) [189]. This system enables the follower vehicle to dynamically adapt and maintain appropriate inter-vehicular distances. Notably, both these studies rely on V2V communication to achieve effective car following behavior. In contrast, [95] introduces a new perspective by realizing multitasking driving, encompassing car following and lane keeping, solely relying on visual sensors without the need for communication with other vehicles. In this approach, a pattern of circles is affixed behind the leading Duckiebot, aiding the ego Duckiebot in sensing distance and speed to mimic human driver behavior. A modular system is deployed, where the perception module extracts compact affordance information, and a DRL algorithm controls the Duckiebot based on this information. Despite its simplicity, this system yields promising results. As an imperative area for advancement, future research should emphasize car following behavior with small-scale cars, aiming to establish a foundational basis for real-world driving applications.
Figure 9: Framwork proposed in [95], the perception module leverages camera images to produce impact attributes regarding the environment, then the DRL control module utilizes the information to control the agent perform car following behavior.
III-D2 Overtaking
Overtaking refers to the maneuver where an autonomous vehicle changes lanes or positions itself to safely pass another vehicle traveling in the same direction on the road. The purpose of overtaking is to move ahead of the slower vehicle safely and efficiently while maintaining proper traffic flow. For scaled cars, it is a challenge to obtain the overtaking maneuver, since during the whole process, a precise localization system is required to avoid collision with the slower vehicle. Therefore, there are just a few papers focusing on the overtaking task with small-scale cars. A first contribution comes from [94], where authors employ a modular system equipped solely with onboard sensors to execute diverse driving tasks, encompassing lane keeping and overtaking for Duckiebots. The framework leverages traditional machine vision technologies to acquire compact affordance information, as discussed in [7], and utilizes LiDAR sensors for distance estimation relative to other vehicles. Afterward, the Long Short-Term Memory Soft Actor-Critic (LSTM-SAC) algorithm assumes the role of the controller. Notably, LSTM aids the agent in recognizing distinct phases within the overtaking process. Demonstrated results affirm the efficacy of this proposed framework, managing both lane keeping and overtaking tasks, and showcasing superior performance when benchmarked against baselines. Another work regarding the overtaking task is [190], where a dual control approach with MPC towards active uncertainty reduction is proposed. This approach automatically balances the exploration-exploitation trade-off, enabling the MuSHR car to actively minimize uncertainty concerning the hidden states of other agents without compromising expected planning performance. The difference compared with [94] from the hardware side is that the vehicle being overtaken will yield, and the usage of a known grid map of the track. The exploration of overtaking tasks remains relatively limited within current research, signaling a crucial need for further investigation and focus in the future.
III-E Obstacle Detection and Avoidance

### v2

@@ IV-A1 Localization and Mapping
Localization and mapping are foundational tasks in autonomous systems, often requiring high precision and real-time performance. A primary consideration in localization is GPS, which offers direct positional data for the vehicle. However, the accuracy of GPS is influenced by various factors, such as atmospheric conditions, satellite geometry, signal blockage (buildings, trees, etc.), multipath interference, and the quality of the GPS receiver. Differential GPS (DGPS) systems and augmentation techniques can be used to enhance GPS accuracy by correcting some of these error sources, but cannot fulfill the accuracy which is extremely important when it is small-scale cars. IMUs, another commonly used sensor, provide high-rate measurements of acceleration and angular velocity across three dimensions but suffer from issues like sensor drift and integration errors, leading to long-term inaccuracies. To address the issues, the two sensors are often integrated through sensor fusion methods, such as Kalman filter (KF) [98] to enhance accuracy and reliability in determining the position and orientation. As in the studies by [119, 120, 121], that use BARC small-scale cars, KF makes the fusion of IMU, wheel encoders, and indoor GPS measurements to achieve an accurate localization. Nevertheless, KFs are limited by assumptions of linearity and Gaussian noise, making them less effective in complex or nonlinear environments. Therefore, [11] and [50] use factor graphs combined with incremental smoothing and mapping 2 (iSAM2) [122], to fuse GPS and IMU measurements and output smoothing estimation. For the issues with non-linear dynamics and non-linear measurement models, [54, 53, 52] use the particle filter to fuse the camera images with IMU, and GPS. [123] also presents an approach to use KF to fuse multiple-camera images with the AprilTag system and wheel odometry, validated using the OptiTrack Motion Capture System, which offers high accuracy but at a significantly higher cost, limiting its practicality for small-scale platforms. Additionally, a single camera can serve as a perception sensor. For example, in [55], CNN is used with a single monocular camera to predict the cost map of the track in front of the vehicle which is directly useable for online trajectory optimization with MPC. Similarly, [124] uses a top-down lane cost map CNN and the YOLOv2 CNN to extract feature-input values and a two-point visual driver control model (TPVDCM) as the controller to control the vehicle. However, these camera-based perception approaches rely solely on visual information, neglecting spatial information, which can lead to limitations in accuracy and robustness under certain conditions, where the camera is not robust enough.
@@
LiDAR-based systems have emerged as a compelling alternative, leveraging 2D LiDAR sensors for localization and mapping via SLAM techniques. A well-established method for localizing robot cars with LiDAR sensors on the pre-defined map involves the use of a particle filter (PF) as explored in [44] with F1TENTH. However, the computational demands of the PF present challenges, especially for computation resources and space-constrained small-scale robot cars. To enhance performance, [125] introduces the Compressed Directional Distance Transform (CDDT) on the RACECAR platform. This method focuses on expediting ray casting within 2D occupancy grid maps and accelerating sensor model computations to mitigate computational expenses. While comprehensive survey papers delve into various SLAM techniques [126, 127, 128, 129], encompassing feature-based, LiDAR-based, visual, and graph SLAM, this study offers a concise overview of three prevalent SLAM methods employed with small-scale cars: GMapping [130, 131, 132], HectorSLAM [133, 134, 135, 136, 137], and Cartographer [138, 139, 140]. While 3D LiDAR remains underutilized due to its size and computational demands, advancements in hardware miniaturization are making 3D LiDAR SLAM techniques like LOAM [141] and LIO-SAM [142] increasingly feasible for small-scale platforms.
@@
Another compelling approach gaining traction involves leveraging ML methods to address mapping and localization challenges. One notable instance is the GALNet proposed in [143] implemented on the Autominy platform. GALNet employs a Deep Neural Network (DNN), utilizing two timestamps of inertial, kinematic, and wheel velocity data to estimate poses effectively. With the success of Transformer architecture in various research fields, [144] introduces the Perception-Action Causal Transformer (PACT) architecture. This model constructs a representation from sensor data by autonomously predicting states and actions over time, laying the groundwork for subsequent task-specific networks for localization and mapping. Pre-trained models are employed initially and later fine-tuned for specific tasks, demonstrating validation on MuSHR with LiDAR sensor data. For image-based localization, [145] proposes an effective framework leveraging a pre-trained local feature transformer (LoFTR). This framework employs a constrained 3D projective transformation between consecutive key images to establish a visual map on Jetbot. These studies underscore the growing interest and potential of ML methodologies, including neural networks and transformer architectures, in revolutionizing mapping and localization paradigms. However, ML approaches are often data-intensive and require extensive pre-training and fine-tuning to generalize effectively. Future research should focus on improving the robustness and adaptability of these methods, integrating advanced sensing technologies, and optimizing computational efficiency to bridge the gap between traditional and data-driven localization paradigms.
TABLE IV: Comparison of methods used for localization and mapping.
Method Sensor Analysis Reference
Kalman Filter GPS; IMU; Encoder $\sqrt{}$ Accuracy improvement $\times$ Asynchrony, Latency $\times$ Noise Assumptions [119, 120, 121, 146]: BARC; [123]: QCar
Particle Filter GPS; IMU; LiDAR $\sqrt{}$ Nonlinear dynamics $\times$ High computational demand [44]: F1TENTH; [125]: RACECAR
GMapping LiDAR $\sqrt{}$ Robust in static environments $\times$ High memory usage [133]: Roborace; [134, 136]: -
HectorSLAM LiDAR $\sqrt{}$ Fast and lightweight $\times$ Performance degrades with noise [130, 131, 132]: -
Cartographer LiDAR $\sqrt{}$ Accuracy, dynamic environments $\times$ High computational demand [147]: F1TENTH; [139]: Innopolis UGV; [138, 140]: -
CNN Camera $\sqrt{}$ Camera only $\times$ FoV of Camera $\times$ No Spatial infomation [55, 124]: AutoRally
GALNet Odometry $\sqrt{}$ Fast $\times$ Dynamic-dependent [143]: Autominy
LoFTR Camera $\sqrt{}$ Camera only $\times$ High computational demand [145]: Jetbot
IV-A2 Lane Recognition
Lane recognition or lane detection is another important task for the perception module. It is the process of identifying and recognizing the lane markings on a road using algorithms. It typically involves detecting lines or curves that represent the boundaries of lanes on the road. For small-scale cars, we categorize the lane detection methods into traditional and ML-based methods.
@@ IV-A2 Lane Recognition
Traditional lane detection methods rely on computer vision techniques to process sensor data, usually captured by cameras, to identify and track lane markings. Feature extraction techniques are then used to compute metrics such as lateral and orientation deviation, which indicate the position of the car relative to the lane. In [9], a multi-step image processing pipeline is employed for lane detection in Duckietown. This pipeline uses techniques such as k-means clustering, the Canny filter [148], HSV colorspace thresholding, and the probabilistic Hough transform [149] to extract lane line segments. These features are further processed by a nonlinear non-parametric histogram filter to estimate the lateral displacement and angular offset relative to the right lane center. Similarly, [150] applies an HSV color-based approach with Gaussian kernel filtering and Sobel edge detection [151], followed by a probabilistic Hough transform for lane detection on the Donkeycar platform. While these traditional methods are computationally efficient and effective for structured environments, their reliance on explicit feature extraction limits their generalization to complex or unstructured scenarios, making them less robust in real-world applications.
@@
ML-based approaches on the other hand improve lane detection by leveraging the predictive capabilities of advanced algorithms. For example, [152] employs a Convolutional Neural Network (CNN) to predict lateral displacement and angular offset. Additionally, [42] introduces a modular system trained on a mix of simulation and real-world datasets, mapping sensor inputs into a shared latent space. Recent advancements in Vision Transformers (ViT) have also been applied to lane detection. In [153], a pre-trained ViT is fine-tuned on limited driving datasets to perform environmental segmentation.
IV-A3 Obstacle Recognition